# Monadic reconstruction of unitary Drinfeld centers and Factorization Homology

### v1

@@ 1. Introduction
Specially for applications in mathematical-physics, the literature concentrates mostly on fusion categories, finite non-semisimple tensor categories, either rigid or of Grothendieck-Verdier type. However, the theory of compact quantum groups ([MR3204665]) and the theory of subfactors ([MR0696688, MR1622812, MR3948170, Hataishi_Palomares_2025])gives rich sources of tensor categories with unitary structure and infinite isomorphism classes of simple objects. They are also rigid, which together with unitarity implies semi-simplicity, and have simple units. Such categories are called unitary tensor categories (UTCs).
@@
where $\Sigma$ is a compact surface. Geometrically, $\mathcal{R}_{K_{\mathbb{C}}}(\Sigma)$ is the moduli space of flat principal $K_{\mathbb{C}}$-bundles on $\Sigma$, trivialized on a point, and the action of $K_{\mathbb{C}}$ corresponds to change of trivialization. Though we can not yet prove this quantization result, we show here that factorization homology can be expressed in terms of $DK_{q}$-C-algebras (Theorem B).
@@
In the fusion case, Müger’s works [MR1966524, MR1966525] assert that the Drinfeld center of a fusion category $\mathcal{C}$ is related to $\mathcal{C}^{mp}\boxtimes\mathcal{C}$ through a Morita context, where $\mathcal{C}^{mp}$ is the category $\mathcal{C}$ with opposition monoidal structure, and $\boxtimes$ is the tensor product of linear categories. Translated to the monadic language of [MR2355605, MR2793022, MR2869176, MR3079759], this tantamonts to the existence of an algebra object $\mathcal{S}$ in $\mathcal{C}^{mp}\boxtimes\mathcal{C}$ such that
@@
Thanks to [MR3007088, MR3431668], it follows that the functor $\operatorname{Disk}_{2}\to\operatorname{C*Cat}$ defined by a C-tensor category $\mathcal{Z}$ with a unitary braiding extends to a symmetric monoidal functor
@@
from the category of compact surfaces, seen as symmetric monoidal (2,1)-category similarly to $\operatorname{Disk}_{2}$. This functor is called factorization homology with coefficients in $\mathcal{Z}$ ([MR3431668]), and satisfies an excision property in terms of collar-gluing of surfaces.
We use Theorem A to study $\int_{(-)}\mathcal{Z}\mathrm{Hilb}(\mathcal{C})$, factorization homology with coefficients in the unitary Drinfeld center of a UTC $\mathcal{C}$. FH naturally leads to the study of pointed module C-categories for unitary Drinfeld centers (a pointed module category is a module category with a distinguished object). Using Theorem A, the following can be proven.
@@
of $\mathcal{C}^{\boxtimes n}$-bimodule C-categories, where the target is the category of left Hilbert $(B_{\Sigma},B_{F}^{\otimes n})$-C-correspondences. This construction is functorial in $\Sigma$. Moreover, the modular group $\Gamma(\Sigma)$ of $\Sigma$ relative to the boundary acts on $B_{\Sigma}$ by *-automorphims, and $B_{F}^{\otimes n}$ is globally fixed under this action.
@@ 1.1. Acknowledgements
I would like to thank Makoto Yamashita, Sergey Neshveyev, Corey Jones, David Jordan, Christian Voigt and Thomas Wasserman for helpful conversations on topics related to this paper. L.H. is supported by the Engineering and Physical Sciences Research Council (EP/X026647/1).
@@ 2.1. Notation
For notational simplicity, when $a$ and $b$ are objects of a tensor category, we may sometimes write $ab$ for their tensor product $a\otimes b$.
@@ A C-category $\mathcal{C}$ consists of
  • A category $\mathcal{C}$ enriched over the category of Banach spaces and contractions;
@@
The following notation is used for the duality functor in rigid tensor categories. On objects, we write $a\mapsto\bar{a}$. On morphisms, we write $(f:a\to b)\mapsto(f^{\vee}:\bar{b}\to\bar{a})$.
Unit objects in tensor categories will generically be denoted by $\mathds{1}$}.
@@
Unitary tensor categories are automatically semisimple and spherical. To simplify the notation, we work only with strictfied unitary tensor categories: the tensor structure is strict and the double dual functor is the identity, $(\bar{\bar{(-)}},(-)^{\vee\vee})=\mathrm{id}$. If a given unitary tensor category is not strictfied on the nose, we replace it by its strictification without mentioning.
@@
Let $\mathcal{C}$ be a unitary tensor category. The algebraic ind-completion $\operatorname{Vec}(\mathcal{C})$ of $\mathcal{C}$ is the category of functors $\mathcal{C}^{\text{op}}\to\operatorname{Vec}$. Morphisms in this category are linear natural transformations.
To specify an object $V\in\operatorname{Vec}(\mathcal{C})$, it suffices, up to isomorphism, to specify its values, or fibers, at the objects in $\operatorname{Irr}(\mathcal{C})$, i.e., a family $\{V(a)\}_{a\in\operatorname{Irr}(\mathcal{C})}$ of vector spaces indexed by $\operatorname{Irr}(\mathcal{C})$. Another useful perspective is that of understanding the object $V=\{V_{a}\}_{a\in\operatorname{Irr}(\mathcal{C})}\in\operatorname{Vec}(\mathcal{C})$ as a formal direct sum
@@
of objects in $\mathcal{C}$, where the irreducible object $a$ appears $\dim V(a)$ times as a direct summand. We say that $V(a)$ is the multiplicity space of $a$ in $V$. Moreover, given $V_{1}=\{V_{1}(a)\}_{a}$ and $V_{2}=\{V_{2}(a)\}_{X}$ in $\operatorname{Vec}(\mathcal{C})$, it follows that
@@
  • objects: functors $\mathcal{C}^{\text{op}}\to\mathrm{Hilb}$;
@@
where on the $(b,c)$-direct summand the inner-product is scaled by a factor of $(d_{b}d_{c})^{-1}$. The purpose of this renormalization is to have a clean interpretation of the induced inner product by means of graphical calculus (see [MR3687214], Section 2).
@@
  • involution: $j^{\mathbb{A}}_{\bar{a}}\circ j^{\mathbb{A}}_{X}=\mathrm{id}_{\mathbb{A}(X)}$, under the identification $\bar{\bar{a}}=a$;
  • unitality: $j^{\mathbb{A}}_{1_{\mathcal{C}}}=\mathrm{id}_{\mathbb{A}(1_{\mathcal{C}})}$, under the identification $\bar{{\mathds{1}}}\simeq{\mathds{1}}$;
@@
If $\mathbb{A}=(\mathbb{A},\mu^{\mathbb{A}})$ is an algebra object in $\operatorname{Vec}(\mathcal{C})$, then the vector spaces $\mathbb{A}(\bar{a}\boxtimes a)$ become algebras, with multiplication given by
$\mathbb{A}(\bar{a}\otimes a)\odot\mathbb{A}(\bar{a}\otimes a)\overset{\mu^{\mathbb{A}}}{\to}\mathbb{A}(\bar{a}\otimes a\otimes\bar{a}\otimes a)\overset{\mathbb{A}(\mathrm{id}_{\bar{a}}{\otimes\bar{R}_{a}^{*}\otimes\mathrm{id}_{a})}}{\to}\mathbb{A}(\bar{a}\boxtimes a).$
@@
Let $\mathbb{A}=(\mathbb{A},\mu^{\mathbb{A}},j^{\mathbb{A}})$ be a *-algebra in $\operatorname{Vec}(\mathcal{C})$. We say that $\mathbb{A}$ is a C-algebra object if, for each object $a$, the *-algebra $\mathbb{A}(\bar{a}\otimes a)$ is a C-algebra.
@@
has a right $\mathcal{C}$-module structure. $\mathbb{A}$ is a C-algebra object if and only if $\mathcal{M}_{\mathbb{A}}$ is a $\mathcal{C}$-module C-category [MR3687214, Definition 25]. In particular, if $\mathbb{A}$ is a C-algebra object, $\mathbb{A}(a)$ has a Banach space structure for every $a\in\mathcal{C}$.
@@ 3. Monadicity of the center
This section enhances the results of [MR2355605, MR2793022, MR2869176, MR3079759], summarized in [MR3674995], in the direction of giving a monadic characterization of $\mathcal{Z}\operatorname{Vec}(\mathcal{C})$ when $\mathcal{C}$ is a unitary tensor category. The strategy and even the techniques used here are recicled from those references, the difference being that at some moments extra care is needed when dealing with ind-objects.
@@ 3.1. Centralizer endofunctors
$X(U\odot(-))\simeq\bar{U}\odot X$
@@
$\mathcal{C}\ni U\mapsto Z_{\operatorname{Vec}}(X)(U)\simeq\bigoplus_{a\in\operatorname{Irr}(\mathcal{C})}X(a\odot U\odot\bar{a}).$
@@
$Z_{\mathrm{Hilb}}(X)=\bigoplus_{a\in\operatorname{Irr}(\mathcal{C})}^{\ell^{2}}\bar{a}\otimes X\otimes a$
@@
$\mathcal{C}\ni U\mapsto Z_{\mathrm{Hilb}}(X)(U)\simeq\bigoplus_{a\in\operatorname{Irr}(\mathcal{C})}^{\ell^{2}}X(a\otimes U\otimes\bar{a}).$
@@
$\rho_{X,U}:=\sum_{a\in\operatorname{Irr}(\mathcal{C})}\sum_{\omega\in\operatorname{ONB}(a,U)}((\omega^{\vee})^{*}\odot\mathrm{id}_{X}\odot\omega^{*})$
@@
$\rho_{X,U_{1}}\circ C_{X}(f^{\text{op}},\mathrm{id}_{U_{1}})$ $=\sum_{a\in\operatorname{Irr}(\mathcal{C})}\sum_{w\in\operatorname{ONB}(a,U_{1})}((\omega^{\vee})^{*}\otimes\mathrm{id}_{X}\otimes\omega^{*})(f^{\vee}\otimes\mathrm{id}_{X}\otimes\mathrm{id}_{U_{1}})$
$=\sum_{a\in\operatorname{Irr}(\mathcal{C})}\sum_{w\in\operatorname{ONB}(a,U_{1})}((f\circ\omega^{*})^{\vee}\otimes\mathrm{id}_{X}\otimes\omega^{*})$
$=\sum_{a\in\operatorname{Irr}(\mathcal{C})}\sum_{v\in\operatorname{ONB}(a,U_{2})}((v^{*})^{\vee}\otimes\mathrm{id}_{X}\otimes(f\circ v^{*}))$
@@
$\rho_{X,U}^{*}\rho_{X,U}$ $=\sum_{i,j\in\operatorname{Irr}(\mathcal{C})}\sum_{\omega_{i}\in\operatorname{ONB}(U_{i},U)}\sum_{\omega_{j}\in\operatorname{ONB}(U_{j},U)}\omega_{j}^{\vee*}\omega_{i}^{\vee}\otimes\mathrm{id}_{X}\otimes\omega_{j}\omega_{i}^{*}$
$=\sum_{i\in\operatorname{Irr}(\mathcal{C})}\sum_{\omega\in\operatorname{ONB}(U_{i},U)}(\omega\omega^{*})^{\vee}\otimes\mathrm{id}_{X}\otimes\omega\omega^{*}$
@@ 3.2. Free half-braidings
We shall now discuss how to enhance the functor $Z_{\mathrm{Hilb}}\in\operatorname{End}(\mathrm{Hilb}(\mathcal{C}))$ to a functor $\mathrm{Hilb}(\mathcal{C})\to\mathcal{Z}_{*}(\mathrm{Hilb}(\mathcal{C}))$. We start by introducing the free half-braidings.
@@
$(\sigma^{X}_{U})_{a,b}:=\left(\frac{d_{a}}{d_{b}}\right)^{1/2}\sum_{\omega\in\operatorname{ONB}(b,a\otimes U)}(\mathrm{id}_{U}\otimes\omega^{\vee}\otimes\mathrm{id}_{X}\otimes\omega^{*})(\bar{R}_{U}\otimes\mathrm{id}_{\bar{a}}\otimes\mathrm{id}_{X}\otimes\mathrm{id}_{a}).$
@@
The functor $\mathcal{F}_{\mathrm{Hilb}}:\mathrm{Hilb}(\mathcal{C})\to\mathcal{Z}_{*}(\mathrm{Hilb}(\mathcal{C}))$ given by
@@
Denoting by $\mathcal{U}_{\mathrm{Hilb}}:\mathcal{Z}_{*}(\mathrm{Hilb}(\mathcal{C}))\to\mathrm{Hilb}(\mathcal{C})$ the forgetful functor, we have $Z_{\mathrm{Hilb}}=\mathcal{U}_{\mathrm{Hilb}}\circ\mathcal{F}_{\mathrm{Hilb}}$, and one could think of establishing an adjunction $\mathcal{F}_{\mathrm{Hilb}}\dashv\mathcal{U}_{\mathrm{Hilb}}$ so that $Z_{\mathrm{Hilb}}$ is the monad of this adjunction. The counit of this adjunction will not be bounded, and thus it will not be a morphism in $\mathrm{Hilb}(\mathcal{C})$. The monadic data will be well defined in the algebraic ind-completion, but we have to modify the free-half braiding construction a little.
Let $X\in\operatorname{Vec}(\mathcal{C})$. For each $U\in\mathcal{C}$, let $(\beta^{X}_{U})_{a,b}:\bar{a}\otimes X\otimes a\otimes U\to U\otimes\bar{b}\otimes X\otimes b$ be given by
$(\beta^{X}_{U})_{a,b}:=\sum_{\omega\in\operatorname{ONB}(b,a\otimes U)}(\mathrm{id}_{U}\otimes\omega^{\vee}\otimes\mathrm{id}_{X}\otimes\omega^{*})(\bar{R}_{U}\otimes\mathrm{id}_{\bar{a}}\otimes\mathrm{id}_{X}\otimes\mathrm{id}_{a}).$
@@
that takes $X\in\operatorname{Vec}_{\mathcal{C}}$ to $(Z_{\operatorname{Vec}}(X),\beta^{X})$.

### v2

@@ 1. Introduction
Specially for applications in mathematical-physics, the literature concentrates mostly on fusion categories, finite non-semisimple tensor categories, either rigid or of Grothendieck-Verdier type. However, the theory of compact quantum groups ([MR3204665]) and the theory of subfactors ([MR0696688, MR1622812, MR3948170, Hataishi_Palomares_2025])gives rich sources of tensor categories with unitary structures and infinite isomorphism classes of simple objects. They are also rigid, which together with unitarity implies semi-simplicity, and have simple units. Such categories are called unitary tensor categories (UTCs).
@@
where $\Sigma$ is a compact surface. Geometrically, $\mathcal{R}_{K_{\mathbb{C}}}(\Sigma)$ is the moduli space of flat principal $K_{\mathbb{C}}$-bundles on $\Sigma$, trivialized on a point, and the action of $K_{\mathbb{C}}$ corresponds to changes of trivialization. Though we can not yet prove this quantization result, we show here that factorization homology can be expressed in terms of $DK_{q}$-C-algebras (Theorem B).
@@
In the fusion case, Müger’s works [MR1966524, MR1966525] assert that the Drinfeld center of a fusion category $\mathcal{C}$ is related to $\mathcal{C}^{mp}\boxtimes\mathcal{C}$ through a Morita context, where $\mathcal{C}^{mp}$ is the category $\mathcal{C}$ with opposition monoidal structure, and $\boxtimes$ is the tensor product of linear categories. Translated to the monadic language of [MR2355605, MR2793022, MR2869176, MR3079759], this tantamounts to the existence of an algebra object $\mathcal{S}$ in $\mathcal{C}^{mp}\boxtimes\mathcal{C}$ such that
@@
Thanks to [MR3007088] and [MR3431668], it follows that the functor $\operatorname{Disk}_{2}\to\operatorname{C*Cat}$ defined by a C-tensor category $\mathcal{Z}$ with a unitary braiding extends to a symmetric monoidal functor
@@
from the category of compact surfaces, seen as a symmetric monoidal (2,1)-category in similar fashion as $\operatorname{Disk}_{2}$. This functor is called factorization homology with coefficients in $\mathcal{Z}$ ([MR3431668]), and satisfies an excision property in terms of collar-gluing of surfaces. It naturally leads to the study of pointed module C-categories for unitary Drinfeld centers (a pointed module category is a module category with a distinguished object).
We use Theorem A to study $\int_{(-)}\mathcal{Z}\mathrm{Hilb}(\mathcal{C})$, factorization homology with coefficients in the unitary Drinfeld center of a UTC $\mathcal{C}$. Using Theorem A, the following can be proven.
@@
of $\mathcal{C}^{\boxtimes n}$-bimodule C-categories, where the target is the category of left Hilbert $(B_{\Sigma},B_{F}^{\otimes n})$-C-correspondences. This construction is functorial in $\Sigma$. Moreover, the modular group $\Gamma(\Sigma)$ of $\Sigma$ relative to the boundary acts on $B_{\Sigma}$ by *-automorphims, and $B_{F}^{\otimes n}$ is globally fixed under this action. .
@@ 1.1. Acknowledgements
I would like to thank Makoto Yamashita, Sergey Neshveyev, Corey Jones, David Jordan, Christian Voigt and Thomas Wasserman for helpful conversations on topics related to this paper. L.H. is supported by the Engineering and Physical Sciences Research Council (EP/X026647/1). For the purpose of Open Access, the author has applied a CC BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission.
@@ 2.1. Notation
For notational simplicity, when $a$ and $b$ are objects of a strict tensor category, we may sometimes write $ab$ for their tensor product $a\otimes b$.
@@ A C-category $\mathcal{C}$ consists of
  • A category $\mathcal{C}$ enriched over the category of Banach spaces and linear contractions;
@@
The following notation is used for the duality functor in rigid tensor categories. On objects, we write $a\mapsto\bar{a}$. On morphisms, we write $(f:a\to b)\mapsto(f^{\vee}:\bar{b}\to\bar{a})$. Unit objects in tensor categories will generically be denoted by $\mathds{1}$}.
@@
Unitary tensor categories are automatically semisimple and spherical. To simplify the notation, we work only with strictified unitary tensor categories: the tensor structure is strict and the double dual functor is the identity, $(\bar{\bar{(-)}},(-)^{\vee\vee})=\mathrm{id}$. If a given unitary tensor category is not strict on the nose, we replace it by its strictification without mentioning.
Recall that a functor between linear categories with direct sums is additive if it is linear at the level of both objects and morphisms.
@@
Let $\mathcal{C}$ be a unitary tensor category. The algebraic ind-completion $\operatorname{Vec}(\mathcal{C})$ of $\mathcal{C}$ is the category of additive functors $\mathcal{C}^{\text{op}}\to\operatorname{Vec}$. Morphisms in this category are linear natural transformations.
To specify an object $V\in\operatorname{Vec}(\mathcal{C})$ up to isomorphism, it suffices to specify its values, or fibers, at the objects in $\operatorname{Irr}(\mathcal{C})$, i.e., a family $\{V(a)\}_{a\in\operatorname{Irr}(\mathcal{C})}$ of vector spaces indexed by $\operatorname{Irr}(\mathcal{C})$. Another useful perspective is that of understanding the object $V=\{V_{a}\}_{a\in\operatorname{Irr}(\mathcal{C})}\in\operatorname{Vec}(\mathcal{C})$ as a formal direct sum
@@
of objects in $\mathcal{C}$, where the irreducible object $a$ appears $\dim V(a)$ times as a direct summand. We say that $V(a)$ is the multiplicity space of $a$ in $V$. Moreover, given $V_{1}=\{V_{1}(a)\}_{a}$ and $V_{2}=\{V_{2}(a)\}_{a}$ in $\operatorname{Vec}(\mathcal{C})$, it follows that
@@
  • objects: additive functors $\mathcal{C}^{\text{op}}\to\mathrm{Hilb}$;
@@
where on the $(b,c)$-direct summand the inner-product is scaled by a factor of $(d_{b}d_{c})^{-1}$. The purpose of this renormalization is to have a clear interpretation of the induced inner product by means of graphical calculus (see [MR3687214], Section 2).
If $a\in\operatorname{Irr}(\mathcal{C})$, for any $H\in\mathrm{Hilb}(\mathcal{C})$ the morphism space $\mathrm{Hilb}(\mathcal{C})(a,H)$ is a Hilbert space with inner product $\langle f,g\rangle=f^{*}\circ g$. We fix a choice of an orthonormal basis for it, and denote it by $O(a,H)$.
2.3. C-algebra objects
@@
  • involution: $j^{\mathbb{A}}_{\bar{a}}\circ j^{\mathbb{A}}_{a}=\mathrm{id}_{\mathbb{A}(a)}$, under the identification $\bar{\bar{a}}=a$;
  • unitality: $j^{\mathbb{A}}_{{\mathds{1}}}=\mathrm{id}_{\mathbb{A}({\mathds{1}})}$, under the identification $\bar{{\mathds{1}}}\simeq{\mathds{1}}$;
@@
If $\mathbb{A}=(\mathbb{A},\mu^{\mathbb{A}})$ is an algebra object in $\operatorname{Vec}(\mathcal{C})$, then the vector spaces $\mathbb{A}(\bar{a}\otimes a)$ become algebras, with multiplication given by
$\mathbb{A}(\bar{a}\otimes a)\odot\mathbb{A}(\bar{a}\otimes a)\overset{\mu^{\mathbb{A}}}{\to}\mathbb{A}(\bar{a}\otimes a\otimes\bar{a}\otimes a)\overset{\mathbb{A}(\mathrm{id}_{\bar{a}}{\otimes\bar{R}_{a}^{*}\otimes\mathrm{id}_{a})}}{\to}\mathbb{A}(\bar{a}\otimes a).$
@@
Let $\mathbb{A}=(\mathbb{A},\mu^{\mathbb{A}},j^{\mathbb{A}})$ be a *-algebra in $\operatorname{Vec}(\mathcal{C})$. We say that $\mathbb{A}$ is a C-algebra object if its fibers are naturally equipped with Banach space structures and, for each object $a$, the *-algebra $\mathbb{A}(\bar{a}\otimes a)$ is a C-algebra with respect to the given Banach norm.
@@
has a right $\mathcal{C}$-module structure. $\mathbb{A}$ is a C-algebra object if and only if $\mathcal{M}_{\mathbb{A}}$ is a $\mathcal{C}$-module C-category [MR3687214, Definition 25].
@@ 3. Monadicity of the center
This section enhances the results of [MR2355605, MR2793022, MR2869176, MR3079759], summarized in [MR3674995], in the direction of giving a monadic characterization of $\mathcal{Z}\operatorname{Vec}(\mathcal{C})$ when $\mathcal{C}$ is a unitary tensor category. The strategy and even the techniques used here are recicled from the given references, the difference being that at some moments extra care is needed when dealing with ind-objects.
@@ 3.1. Centralizer endofunctors
$X(U\otimes(-))\simeq\bar{U}\odot X$
@@
Proof.
Indeed, for any $a\in\mathcal{C}$,
$X(U\otimes a)\simeq\operatorname{Vec}(\mathcal{C})(U\otimes a,X)\simeq\operatorname{Vec}(\mathcal{C})(a,\bar{U}\odot X)\simeq(\bar{U}\odot X)(a)\ ,$
where the penultim isomorphism is given by the extension $\Phi^{l}_{U}$ to $\operatorname{Vec}(\mathcal{C})$, which is well defined. ∎
@@
$\mathcal{C}^{\text{op}}\ni U\mapsto Z_{\operatorname{Vec}}(X)(U)\simeq\bigoplus_{a\in\operatorname{Irr}(\mathcal{C})}X(a\odot U\odot\bar{a}).$
For the next definition, observe that if $(X_{i})_{i\in I}$ is an arbitrary family of objects in $\mathrm{Hilb}(\mathcal{C})$, we can definite their Hilbert space object direct sum $\bigoplus_{i\in I}^{\ell^{2}}X_{i}$ as the additive functor $\mathcal{C}^{\text{op}}\to\mathrm{Hilb}$ given by
$\mathcal{C}^{\text{op}}\ni U\mapsto\bigoplus_{i\in I}^{\ell^{2}}X_{i}(U)\ .$
@@
$Z_{\mathrm{Hilb}}(X)=\bigoplus_{a\in\operatorname{Irr}(\mathcal{C})}^{\ell^{2}}\bar{a}\otimes X\otimes a\ ,$
@@
$\mathcal{C}^{\text{op}}\ni U\mapsto Z_{\mathrm{Hilb}}(X)(U)\simeq\bigoplus_{a\in\operatorname{Irr}(\mathcal{C})}^{\ell^{2}}X(a\otimes U\otimes\bar{a}).$
@@
$\rho_{X,U}:=\sum_{a\in\operatorname{Irr}(\mathcal{C})}\sum_{\omega\in O(a,U)}((\omega^{\vee})^{*}\odot\mathrm{id}_{X}\odot\omega^{*})$
@@
$\rho_{X,U_{1}}\circ C_{X}(f^{\text{op}},\mathrm{id}_{U_{1}})$ $=\sum_{a\in\operatorname{Irr}(\mathcal{C})}\sum_{w\in O(a,U_{1})}((\omega^{\vee})^{*}\otimes\mathrm{id}_{X}\otimes\omega^{*})(f^{\vee}\otimes\mathrm{id}_{X}\otimes\mathrm{id}_{U_{1}})$
$=\sum_{a\in\operatorname{Irr}(\mathcal{C})}\sum_{w\in O(a,U_{1})}((f\circ\omega^{*})^{\vee}\otimes\mathrm{id}_{X}\otimes\omega^{*})$
$=\sum_{a\in\operatorname{Irr}(\mathcal{C})}\sum_{v\in O(a,U_{2})}((v^{*})^{\vee}\otimes\mathrm{id}_{X}\otimes(f\circ v^{*}))$
@@
$\rho_{X,U}^{*}\rho_{X,U}$ $=\sum_{i,j\in\operatorname{Irr}(\mathcal{C})}\sum_{\omega_{i}\in O(U_{i},U)}\sum_{\omega_{j}\in O(U_{j},U)}\omega_{j}^{\vee*}\omega_{i}^{\vee}\otimes\mathrm{id}_{X}\otimes\omega_{j}\omega_{i}^{*}$
$=\sum_{i\in\operatorname{Irr}(\mathcal{C})}\sum_{\omega\in O(U_{i},U)}(\omega\omega^{*})^{\vee}\otimes\mathrm{id}_{X}\otimes\omega\omega^{*}$
@@ 3.2. Free half-braidings
We shall now discuss how to enhance the functor $Z_{\mathrm{Hilb}}\in\operatorname{End}(\mathrm{Hilb}(\mathcal{C}))$ to a functor $\mathrm{Hilb}(\mathcal{C})\to\mathcal{Z}\mathrm{Hilb}(\mathcal{C})$. We start by introducing the free half-braidings.
@@
$(\sigma^{X}_{U})_{a,b}:=\left(\frac{d_{a}}{d_{b}}\right)^{1/2}\sum_{\omega\in O(b,a\otimes U)}(\mathrm{id}_{U}\otimes\omega^{\vee}\otimes\mathrm{id}_{X}\otimes\omega^{*})(\bar{R}_{U}\otimes\mathrm{id}_{\bar{a}}\otimes\mathrm{id}_{X}\otimes\mathrm{id}_{a}).$
@@
The functor $\mathcal{F}_{\mathrm{Hilb}}:\mathrm{Hilb}(\mathcal{C})\to\mathcal{Z}\mathrm{Hilb}(\mathcal{C})$ given by
@@
Denoting by $\mathcal{U}_{\mathrm{Hilb}}:\mathcal{Z}\mathrm{Hilb}(\mathcal{C})\to\mathrm{Hilb}(\mathcal{C})$ the forgetful functor, we have $Z_{\mathrm{Hilb}}=\mathcal{U}_{\mathrm{Hilb}}\circ\mathcal{F}_{\mathrm{Hilb}}$, and one could think of establishing an adjunction $\mathcal{F}_{\mathrm{Hilb}}\dashv\mathcal{U}_{\mathrm{Hilb}}$ so that $Z_{\mathrm{Hilb}}$ is the monad of this adjunction. The counit of this adjunction will not be bounded, i.e. not a morphism in $\mathrm{Hilb}(\mathcal{C})$. The monadic data will be well defined in the algebraic ind-completion, but we have to modify the free-half braiding construction a little.
Let $X\in\operatorname{Vec}(\mathcal{C})$. For each $U\in\mathcal{C}$, let $(\beta^{X}_{U})_{a,b}:\bar{a}\odot X\odot a\odot U\to U\odot\bar{b}\odot X\odot b$ be given by
$(\beta^{X}_{U})_{a,b}:=\sum_{\omega\in O(b,a\otimes U)}(\mathrm{id}_{U}\odot\omega^{\vee}\odot\mathrm{id}_{X}\odot\omega^{*})(\bar{R}_{U}\odot\mathrm{id}_{\bar{a}}\odot\mathrm{id}_{X}\odot\mathrm{id}_{a}).$
@@
that takes $X\in\operatorname{Vec}(\mathcal{C})$ to $(Z_{\operatorname{Vec}}(X),\beta^{X})$.